-5 (2×-9)-10×
I multiplied -5 and 2x=10×
I then multiple -5 and -9=45
-5×-10×
20× +45=

Found 2 solutions by richwmiller, ewatrrr:
Answer by richwmiller(17219)

already answered

Answer by ewatrrr(24785)

-10x + 45 = 10x Yes. Good
Work
45 = 20x
45/20 = x = 9/4 0r 2 1/4